If there are 6 x10^9 people on the Earth and there is enough freshwater to provide each person with 8.79 x10^5 L, how much freshwater is on the Earth?
Answer by checkley71(8403)

6*10^9*8.79*10^5=52.74*10^14=5.274*10^15 GALLONS OF WATER.